OpenStudy (anonymous):

Help!!!! I will fan and give medals!!!! Help!!! Find the zeros of the polynomial function. f(x) = x2 + 9x + 20

OpenStudy (phi):

I would try to factor it. step 1: list all pairs of numbers that when you multiply them, you get 20 can you do that ?

OpenStudy (anonymous):

Get a hint from the plot http://www.wolframalpha.com/input/?i=plot++x%5E2+%2B+9x+%2B+20+from+x+%3D+-6+to++x%3D-3

OpenStudy (anonymous):

Notcie that 4+5=9 4(5)=20 Can you factor it now?

OpenStudy (anonymous):

i dont understand

OpenStudy (anonymous):

Notice that \[ (a+x) (b+x)=x^2 +a x+b x+ a b=x^2+ (a+b)x+ a b \]

OpenStudy (anonymous):

Your a=4 and your b=5

OpenStudy (anonymous):

\[ x^2+9 x+20=(x+4) (x+5)=0 \] For which valuee your ploynomial is zero?

OpenStudy (anonymous):

Can you finish it?

OpenStudy (anonymous):

no ive never done this

OpenStudy (anonymous):

The zeros are x=-4 and x=-5. If you replace x by -4 or x by -5, your polynomial would be equal to zero

OpenStudy (anonymous):

A. 2, 10 B. –4, –5 C. –2, –10 D. 4, 5

OpenStudy (anonymous):

\[ (-4)^2 + 9(-4) +20= 16-36 +20=0\\ (-5)^2 + 9(-5) +20=25-45+20=0 \]

OpenStudy (anonymous):

so B.

OpenStudy (anonymous):

B is your answer

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

can you help me with more please

OpenStudy (anonymous):

Post it as new problem and tag me in it

OpenStudy (anonymous):

ok